# Graduate Student Member, IEEE

### v1

@@ 1 Introduction
As mission designs grow more complex and the number of planned lunar missions increases [spacenews_novaspace_2025], the demand for navigation solutions that provide greater onboard autonomy in orbit determination and can scale with future mission needs has become a practical and strategic necessity. Historically, spacecraft operating beyond Earth orbit have depended primarily on ground-based Radio Frequency (RF) tracking networks, such as Deep Space Network (DSN), European Space Tracking Network (ESTRACK), ISRO Telemetry, Tracking and Command Network (ISTRAC), and Chinese Deep Space Network (CDSN) to support navigation, guidance, and maneuvering. In parallel, recent progress in lunar exploration has highlighted the potential of Global Navigation Satellite System (GNSS) as a complementary resource for navigation in the lunar domain, encompassing cis-lunar space and the lunar surface, making the exploitation of these signals an increasingly compelling complement [JournalDiMissione].
@@
In particular, the Lunar GNSS Receiver Experiment (LuGRE) [parker2022lunar, konitzer2024science] — a joint NASA–ASI experiment hosted on Firefly Aerospace Inc.’s Blue Ghost Mission 1 (BGM1) lunar lander under NASA’s Commercial Lunar Payload Services (CLPS) program — was designed to acquire and track Global Positioning System (GPS) and Galileo signals in the L1/E1 and L5/E5a frequency bands along its trajectory to the Moon and during lunar surface operations, thereby exploiting the resulting observables to estimate the onboard receiver state, i.e., computing its Position, Velocity, Time (PVT). Launched in January 2025, LuGRE achieved historic milestones, including multiple GNSS-only PVT fixes along its trajectory and the first GNSS-only position fix on the lunar surface [JournalDiMissione].
@@
Beyond computing onboard PVT solutions, the LuGRE payload [tedesco2023deep, pulliero2023space] collected more than 106 hours of unprecedented raw GNSS observables, as well as approximately $12$ s of In-phase and Quadrature Samples (IQS) snapshots, with snapshot durations ranging from 200 ms to 2 s, collected across different operational windows. As a result, the openly available LuGRE dataset [parker2025lunar] provides, in addition to state-estimation solutions, both IQS snapshots and GNSS receiver observables collected throughout the mission.
@@
This paper leverages the unique IQS snapshots obtained by LuGRE [parker2022lunar, konitzer2024science], to demonstrate the availability of additional navigation signals from systems not originally exploited during mission operations, i.e., Beidou, the Quasi-Zenith Satellite System (QZSS), the Indian Regional Navigation Satellite System (IRNSS) (also known as NavIC) and Satellite-based Augmentation System (SBAS) constellations. A list of the constellations transmitting SBAS signals is reported in Table 1
@@
Beyond the experimental demonstration itself, the information extracted from the openly available IQS snapshots [parker2025lunar] is used to calibrate and validate an advanced extended Space Service Volume (SSV) simulator. This calibration, grounded directly on in-space experimental data, enables realistic simulation of LuGRE operations while explicitly accounting for the contribution of additional GNSS, Regional Navigation Satellite System (RNSS), and SBAS constellations not processed during LuGRE operations. As a result, the simulator provides, for the first time, an experimentally anchored assessment of multi-constellation signal availability in the cis-lunar and lunar environments.
@@
Several experiments have demonstrated the feasibility of using GNSS signals beyond Low Earth Orbit (LEO). Notably, National Aeronautics and Space Administration (NASA)’s Magentospheric Muliscale (MMS) mission exploited GPS signals at distances of up to approximately 50% of the Earth–Moon distance [mms2016navi, winternitz2017new, mmsRecordNews]. Complementarily, the Canadian Advanced Nanosatellite eXperiment-2 (CanX-2) mission demonstrated SBAS-based ranging in LEO [kahr2016analysis, sarda2010canadian], revealing the need for dedicated corrections to mitigate residual clock offsets on the order of 100 m and for ephemerides more accurate than those provided by the SBAS Satellite Vehicles (SVs). The missions discussed above did not lead to the release of publicly available datasets in the form of signal samples. The LuGRE mission, instead, provides openly accessible data [parker2025lunar]. Accordingly, the following section presents an overview of LuGRE, with particular emphasis on the LuGRE payload and the IQS data collection operations.
@@ 1.1 The LuGRE Payload
The payload’s HGA was specifically designed for the reception of GNSS signals in the L1/E1 and L5/E5a bands. It provides a peak boresight gain of 15.35 dBic and includes an integrated notch filter to suppress out-of-band interference. During the transit phase, the antenna, mounted on a gimballed platform, maintained Earth-pointing by adjusting the spacecraft attitude while keeping the gimbal fixed, resulting in an Earth-Off-Pointing angle consistently below the nominal limit of 1°. During the surface phase, the active gimbal steering further improved alignment, achieving Earth-Off-Pointing angles better than 0.21° [JournalDiMissione]. Consequently, antenna pointing error can be neglected for the analysis conducted in this work.
The LuGRE payload was integrated into the BGM1 lander as a scientific instrument and operated in full compliance with strict do-no-harm requirements [JournalDiMissione, pulliero2023space]. The lander did not rely on the payload for any spacecraft maneuvers, as the receiver was not interfaced with the Guidance Navigation and Control (GNC) system.
@@
The IQS snapshots collected in SC mode were sampled at frequencies between 4 MHz and 24 MHz, with 4-bit or 8-bit quantization. Due to the limited downlink bandwidth shared among all payloads and subsystems, sampling and quantization choices directly impacted the available snapshot duration, limiting each snapshot to a maximum of two seconds, and higher-fidelity sampling configurations could not be used [nardin2023tracking]. Although the SC and RTP modes were mutually exclusive, the two modes were frequently executed in succession within the same operational window. Detailed timing, sequencing, and duration of these mode transitions for each operation are documented in the LuGRE operations summary table (OPTABLE) contained in the mission’s public dataset [parker2025lunar] and in [JournalDiMissione].
@@ 1.2 LuGRE Sample Capture Operations
The start time of each SC operation was selected to maximize the availability of GPS and Galileo signals, when possible. Satellite visibility forecasts used during mission planning were generated using the Goddard Enhanced Onboard Navigation System (GEONS) and GEONS Ground Matlab Simulation (GGMS) tools [JournalDiMissione].
The present article examines only data originated by operations featuring a SC, and all results are obtained from the post-processing of the corresponding IQS snapshots. Table 2 summarizes these operations, indicating, for each one, the ADC configuration and the duration of the acquired signal snapshots. When both L1/E1 and L5/E5a IQS snapshots were collected, the reported duration refers to both bands. Only operations in which the receiver was configured in SC mode are listed. The table also provides the Coordinated Universal Time (UTC) start time of each SC operation, denoted as SCstartSC\ start, together with the associated Julian day ($DoY$). UTC timestamps assigned by the onboard computer to SC operations, along with additional receiver status information, are available in the publicly released dataset [parker2025lunar].
Fig. 1 shows an example of the estimated Power Spectral Density (PSD) of the IQS snapshots for one of the possible ADC configurations on L1 band. As shown in [JournalDiMissione], the spurious frequencies observed are not expected to affect the measurement quality discussed in this work. Nonetheless, all operations exhibit a significant frequency drift that is inconsistent with the expected receiver dynamics. For the sake of completeness, Fig. 2 presents the frequency drift relative to the start time of the SC operation for a representative set of cases, where the frequency offset is estimated over multiple IQS snapshot windows. The frequency offset drift is expected to affect the acquisition stage inducing a reduced performance by introducing a correlation loss dependent on the acquisition stage parameters. However, the RTP dataset is less affected by the drift [JournalDiMissione], it is likely that the drift is introduced by the receiver clock when a reset of the receiver is performed, this happens when selecting the SC and at the RTP mode as well; anyway in RTP mode the receiver clock requires many seconds between the acquisition and the beginning of the signal tracking to stabilize itself. A further analysis on the root causes is out of scope and left for future works.
@@ 2.1 Signal Acquisition Architecture and Methods
The adopted signal acquisition method is based on a hybrid coherent and non-coherent integration scheme with code Doppler compensation, as investigated in previous analyses [nardin2022acquisition]. For background on standard GNSS signal acquisition architectures and methods, the interested reader is referred to [kaplan2017understanding]. The set of acquisition parameters chosen for the acquisition of the GNSS, RNSS and SBAS signals under analysis is provided in Table 3. As reported in the literature [van1991new, tsui2004fundamentals, borio2008statistical], signal acquisition can be performed efficiently using a Discrete Fourier Transform (DFT)-based approach, which significantly reduces computational complexity. The mathematical formulation of the process is provided in the following, while a schematic representation of the Cross Ambiguity Function (CAF) computation and of the Carrier to Noise Density Ratio ($C/N_{0}$) estimation process is reported in Fig. 3 [kaplan2017understanding]. Given the high dynamics of the scenario and the behavior of the payload clocks during operations, the Doppler effect and the receiver clock drift significantly affect the correlation between the received and locally generated codes [foucras2014detailed] [sciacca2025acquiring]. As a result, the code Doppler must be accounted for when generating the local carrier and the local code.
@@ 2.2 Acquisition threshold
where $\chi^{-2}$ is the inverse chi-square function. The term $\hat{\sigma}^{2}_{n}$ refers to the estimated noise variance measured at the acquisition stage. To obtain $\hat{\sigma}^{2}_{n}$, according to the literature [betz2001effect, gomez2018c], it is possible to consider the correlation under the null hypothesis $H_{0}$ in the Neyman-Pearson sense [neyman1933ix] as the portion of CAF $S_{H_{0}}$ located away from the correlation peak of at least one chip in the code phase domain. The expression of the noise variance estimate $\hat{\sigma}^{2}_{n}$ is
@@
2.3 C/No characterization
Several methods for estimating $C/N_{0}$ are available in the literature; some rely on the outputs of the correlators at the tracking stage [pini2008performance], such as the Moment Method (MM) and Narrow-Wideband Power Ratio (NWPR) approaches. However, given the reduced length of the IQS snapshots, it is relevant to reliably estimate $C/N_{0}$ during the acquisition stage, for this purpose, the estimator chosen in this analysis is based on the Non coherent Post Detection Integration (NPDI) estimator introduced in [gomez2018c]. The estimate of $C/N_{0}$, in dB-Hz, is obtained as
@@ 2.3 C/No characterization
This estimator is specifically intended for non coherent integration and it is shown hereafter that it has a reduced bias in the range of $C/N_{0}$ values considered in this work.
To understand how the bias introduced by the estimator affects the measurements, a set of simulations is conducted to evaluate the range in which the estimator is in its unbiased regime. A synthetic GPS L1 C/A signal is generated with a sampling frequency of $8$ MHz and a quantization depth of 8 bits. Fifteen $C/N_{0}$ values, linearly spaced between 15 and 55 dB-Hz, are tested. For each $C/N_{0}$ level, 90 trials with independent noise realizations are evaluated using the estimator defined in (10). The number of non-coherent integrations is set to $K=15$, with a coherent integration time $T_{\text{coh}}=4\ \text{ms}$. As shown in Fig. 4, the NPDI based estimator exhibits a bias at higher simulated $C/N_{0}$ values, causing the estimated value to negatively diverge from the true one. A similar behavior is also observed at lower $C/N_{0}$ levels where estimated values become quickly unreliable as $C/N_{0}$ drops below 25 dB-Hz. Taking into account the chosen acquisition parameters and the simulation setup, the estimated $C/N_{0}$ range between 25 and approximately 40 dB-Hz are assumed reliable. Given the limited number of non-coherent integrations and the length of the coherent integration times adopted in this work, the $C/N_{0}$ values that can be observed considered the receiver sensitivity are expected to fall within this range in most of the tested scenarios.
@@
Figure 4: Comparison between expected and measured $C/N_{0}$. Generated $C/N_{0}$ (green), estimated $C/N_{0}$ (red). Coherent integration time $T_{coh}:\ 4\ ms$, non coherent integrations: K: 15K:\ 15.
@@ 2.4 Extended Space Service Volume Simulator
The extended SSV simulator used in this work models the GNSS, RNSS, and SBAS constellations relevant to the analysis. It propagates broadcast ephemerides to determine satellite positions and generates the corresponding observables, including $C/N_{0}$, geometric Doppler shift, and pseudoranges. To estimate satellite visibility and received $C/N_{0}$ for L1 C/A and L5 signals at the spacecraft location, the simulator assigns to each satellite an antenna radiation pattern derived from the available literature. Table 4 summarizes the pattern adopted for each constellation. The extended SSV simulator used in this work models the GNSS, RNSS, and SBAS constellations relevant to the analysis. It propagates broadcast ephemerides to determine satellite positions and generates the corresponding observables, including $C/N_{0}$, geometric Doppler shift, and pseudoranges. To estimate satellite visibility and received $C/N_{0}$ for L1 C/A and L5 signals at the spacecraft location, the simulator assigns to each satellite an antenna radiation pattern derived from the available literature. Table 4 summarizes the pattern adopted for each constellation.
High-fidelity three-dimensional gain patterns are available for GPS, Galileo, and QZSS on both L1 and L5; these are directly incorporated into the simulation. For BeiDou, complete antenna patterns have not yet been published at the time of writing. The simulator therefore employs the Min-Model, which provides a conservative representation of the main lobe for each band [unoosa_icg17_wgb_07] and reconstructs the corresponding pattern through interpolation of the published values.
For other RNSS systems, detailed radiation patterns are similarly unavailable. In the case of GPS-Aided GEO Augmented Navigation (GAGAN) and IRNSS, an approximate antenna pattern is derived from CanX-2 mission measurements reported in [kahr2016analysis]. This pattern is constructed by interpolating the available points conservatively, preserving the main lobe while assigning low gain at large off-boresight angles. For the remaining SBAS constellations, except for GAGAN, the BeiDou Min-Model is used for Geostationary Earth Orbit (GEO) and Inclined Geosynchronous Orbit (IGSO) satellites.
@@
For the remaining GNSS, RNSS, and SBAS constellations, calibration relies exclusively on the $C/N_{0}$ estimates obtained from the IQS snapshots collected during SC operations. Although this provides a less accurate calibration, due to the limited number of available measurements and the absence of detailed antenna patterns, it enables, for the first time, a simulation of multi-constellation signal availability in the cis-lunar environment based directly on experimental observations acquired in cislunar space and at the lunar surface. It is important to note that the power distribution among the various signal components for GPS, as well as for other constellations, may vary over time, for instance, in the case of GPS, due to the feature called flex power mode [steigenberger2019flex].
@@
GPS Official antenna pattern pre-flight measurements released by Lockheed Martin and Boeing. [uscg_navcen_gps_tech_refs]
Galileo GRAP model. [menzione2024reconstruction]
BeiDou (BDS-3) Min-model of the antenna main lobe, available per band and orbit type. [unoosa_icg17_wgb_07] [lin2020bds]
QZSS Official antenna pattern released by the Japan Aerospace Exploration Agency (JAXA). [qzss_antenna_patterns]
NavIC (IRNSS) Pattern available only for a submodule of the antenna. The patterns used correspond to BeiDou BDS-3 (B1) GEO-IGSO. [RajeevGAGAN2005]
@@
** Except for GAGAN [kahr2016analysis].

### v2

@@ 1 Introduction
As mission designs grow more complex and the number of planned lunar missions increases [23], the demand for navigation solutions that provide greater onboard autonomy in orbit determination and can scale with future mission needs has become a practical and strategic necessity. Historically, spacecraft operating beyond Earth orbit have depended primarily on ground-based Radio Frequency (RF) tracking networks, such as Deep Space Network (DSN), European Space Tracking Network (ESTRACK), ISRO Telemetry, Tracking and Command Network (ISTRAC), and Chinese Deep Space Network (CDSN) to support navigation, guidance, and maneuvering. In parallel, recent progress in lunar exploration has highlighted the potential of Global Navigation Satellite System (GNSS) as a complementary resource for navigation in the lunar domain, encompassing cis-lunar space and the lunar surface, making the exploitation of these signals an increasingly compelling complement [17].
@@
In particular, the Lunar GNSS Receiver Experiment (LuGRE) [18, 10] — a joint NASA–ASI experiment hosted on Firefly Aerospace Inc.’s Blue Ghost Mission 1 (BGM1) lunar lander under NASA’s Commercial Lunar Payload Services (CLPS) program — was designed to acquire and track Global Positioning System (GPS) and Galileo signals in the L1/E1 and L5/E5a frequency bands along its trajectory to the Moon and during lunar surface operations, thereby exploiting the resulting observables to estimate the onboard receiver state, i.e., computing its Position, Velocity, Time (PVT). Launched in January 2025, LuGRE achieved historic milestones, including multiple GNSS-only PVT fixes along its trajectory and the first GNSS-only position fix on the lunar surface [17].
@@
Beyond computing onboard PVT solutions, the LuGRE payload [25, 20] collected more than 106 hours of unprecedented raw GNSS observables, as well as approximately $12$ s of In-phase and Quadrature Samples (IQS) snapshots, with snapshot durations ranging from 200 ms to 2 s, collected across different operational windows. As a result, the openly available LuGRE dataset [16] provides, in addition to state-estimation solutions, both IQS snapshots and GNSS receiver observables collected throughout the mission.
@@
This paper leverages the unique IQS snapshots obtained by LuGRE [18, 10], to demonstrate the availability of additional navigation signals from systems not originally exploited during mission operations, i.e., BeiDou, the Quasi-Zenith Satellite System (QZSS), the Indian Regional Navigation Satellite System (IRNSS) (also known as NavIC) and Satellite-based Augmentation System (SBAS) constellations. A list of the constellations transmitting SBAS signals is reported in Table 1.
@@
Beyond the experimental demonstration itself, the information extracted from the openly available IQS snapshots [16] is used to calibrate and validate an advanced extended Space Service Volume (SSV) simulator. This calibration, grounded directly on in-space experimental data, enables realistic simulation of LuGRE operations while explicitly accounting for the contribution of additional GNSS, Regional Navigation Satellite System (RNSS), and SBAS constellations not processed during LuGRE operations. As a result, the simulator provides, for the first time, an experimentally anchored assessment of multi-constellation signal availability in the cis-lunar and lunar environments.
@@
2 Background
Several experiments have demonstrated the feasibility of using GNSS signals beyond Low Earth Orbit (LEO). Notably, National Aeronautics and Space Administration (NASA)’s Magnetospheric Multiscale (MMS) mission exploited GPS signals at distances of up to approximately 50% of the Earth–Moon distance [30, 29, 1]. Complementarily, the Canadian Advanced Nanosatellite eXperiment-2 (CanX-2) mission demonstrated SBAS-based ranging in LEO [8, 21], revealing the need for dedicated corrections to mitigate residual clock offsets on the order of 100 m and for ephemerides more accurate than those provided by the SBAS Satellite Vehicles (SVs). The missions discussed above did not lead to the release of publicly available datasets in the form of signal samples. The LuGRE mission, instead, provides openly accessible data [16]. Accordingly, this section presents an overview of LuGRE, with particular emphasis on the LuGRE payload and the IQS data collection operations.
@@ 2.1 The LuGRE Payload
The payload’s HGA was specifically designed for the reception of GNSS signals in the L1/E1 and L5/E5a bands. It provides a peak boresight gain of 15.35 dBic and includes an integrated notch filter to suppress out-of-band interference. During the transit phase operations, the antenna, mounted on a gimballed platform, was kept pointed toward the Earth by adjusting the spacecraft attitude while keeping the gimbal fixed, resulting in an Earth-Off-Pointing angle consistently below the nominal limit of 1°. During the surface phase, the active gimbal steering further improved alignment, achieving Earth-Off-Pointing angles better than 0.21° [17]. Consequently, antenna pointing error can be neglected for the analysis conducted in this work.
The LuGRE payload was integrated into the BGM1 lander as a scientific instrument and operated in full compliance with strict do-no-harm requirements [17, 20]. The lander did not rely on the payload for any spacecraft maneuvers, as the receiver was not interfaced with the Guidance Navigation and Control (GNC) system.
@@
The IQS snapshots collected in SC mode were sampled at frequencies between 4 MHz and 24 MHz, with 4-bit or 8-bit quantization. Due to the limited downlink bandwidth shared among all payloads and subsystems, sampling and quantization choices directly impacted the available snapshot duration, limiting each snapshot to a maximum of two seconds, and higher-fidelity sampling configurations could not be used [13]. Although the SC and RTP modes were mutually exclusive, the two modes were frequently executed in succession within the same operational window. Detailed timing, sequencing, and duration of these mode transitions for each operation are documented in the LuGRE operations summary table (OPTABLE) contained in the mission’s public dataset [16] and in [17].
@@ 2.2 LuGRE Sample Capture Operations
The start time of each SC operation was selected to maximize the availability of GPS and Galileo signals, when possible. Satellite visibility forecasts used during mission planning were generated using the Goddard Enhanced Onboard Navigation System (GEONS) and GEONS Ground Matlab Simulation (GGMS) tools [17].
The present article examines only data originated by operations featuring a SC, and all results are obtained from the post-processing of the corresponding IQS snapshots. Table 2 summarizes these operations, indicating, for each one, the ADC configuration and the duration of the acquired signal snapshots. When both L1/E1 and L5/E5a IQS snapshots were collected, the reported duration refers to both bands. Only operations in which the receiver was configured in SC mode are listed. The table also provides the Coordinated Universal Time (UTC) start time of each SC operation, denoted as SCstartSC\ start, together with the associated Julian day ($DoY$). UTC timestamps assigned by the onboard computer to SC operations, along with additional receiver status information, are available in the publicly released dataset [16].
Fig. 1 shows an example of the estimated Power Spectral Density (PSD) of the IQS snapshots for one of the possible ADC configurations on L1 band. As shown in [17], the spurious frequencies observed are not expected to affect the measurement quality discussed in this work. Nonetheless, all operations exhibit a significant frequency drift that is inconsistent with the expected receiver dynamics. For the sake of completeness, Fig. 2 presents the frequency drift relative to the start time of the SC operation for a representative set of cases, where the frequency offset is estimated over multiple IQS snapshot windows. The frequency offset drift is expected to affect the acquisition stage inducing a reduced performance by introducing a correlation loss dependent on the acquisition stage parameters. However, the RTP dataset is less affected by the drift [17], it is likely that the drift is introduced by the receiver clock when a reset of the receiver is performed, this happens when selecting the SC and at the RTP mode as well; anyway in RTP mode the receiver clock requires many seconds between the acquisition and the beginning of the signal tracking to stabilize itself. A further analysis on the root causes is out of scope and left for future works.
@@ 3.1 Signal Acquisition Architecture and Methods
The adopted signal acquisition method is based on a hybrid coherent and non-coherent integration scheme with code Doppler compensation, as investigated in previous analyses [14]. For background on standard GNSS signal acquisition architectures and methods, the interested reader is referred to [9]. The set of acquisition parameters chosen for the acquisition of the GNSS, RNSS, and SBAS signals under analysis is provided in Table 3. As reported in the literature [28, 26, 3], signal acquisition can be performed efficiently using a Discrete Fourier Transform (DFT)-based approach, which significantly reduces computational complexity. The mathematical formulation of the process is provided in the following, while a schematic representation of the Cross Ambiguity Function (CAF) computation and of the Carrier to Noise Density Ratio ($C/N_{0}$) estimation process is reported in Fig. 3 [9]. Given the high dynamics of the scenario and the behavior of the payload clocks during operations, the Doppler effect and the receiver clock drift significantly affect the correlation between the received and locally generated codes [4, 22]. As a result, the code Doppler must be accounted for when generating the local carrier and the local code.
@@ 3.2 Acquisition threshold
where $\chi^{-2}$ is the inverse chi-square function. The term $\hat{\sigma}^{2}_{n}$ refers to the estimated noise variance measured at the acquisition stage. To obtain $\hat{\sigma}^{2}_{n}$, according to the literature [2, 5], it is possible to consider the correlation under the null hypothesis $H_{0}$ in the Neyman-Pearson sense [15] as the portion of CAF $S_{H_{0}}$ located away from the correlation peak of at least one chip in the code phase domain. The expression of the noise variance estimate $\hat{\sigma}^{2}_{n}$ is
@@
3.3 $C/N_{0}$ characterization
Several methods for estimating $C/N_{0}$ are available in the literature; some rely on the outputs of the correlators at the tracking stage [19], such as the Moment Method (MM) and Narrow-Wideband Power Ratio (NWPR) approaches. However, given the reduced length of the IQS snapshots, it is relevant to reliably estimate $C/N_{0}$ during the acquisition stage, for this purpose, the estimator chosen in this analysis is based on the Non-coherent Post Detection Integration (NPDI) estimator introduced in [5]. The estimate of $C/N_{0}$, in dB-Hz, is obtained as
@@ 3.3 $C/N_{0}$ characterization
This estimator is specifically intended for non-coherent integration and it is shown hereafter that it has a reduced bias in the range of $C/N_{0}$ values considered in this work.
To understand how the bias introduced by the estimator affects the measurements, a set of simulations is conducted to evaluate the range in which the estimator is in its unbiased regime. A synthetic GPS L1 C/A signal is generated with a sampling frequency of $8$ MHz and a quantization depth of 8 bits. Fifteen $C/N_{0}$ values, linearly spaced between 15 and 55 dB-Hz, are tested. For each $C/N_{0}$ level, 90 trials with independent noise realizations are evaluated using the estimator defined in (10). The number of non-coherent integrations is set to $K=15$, with a coherent integration time $T_{\text{coh}}=4\ \text{ms}$. As shown in Fig. 4, the NPDI based estimator exhibits a bias at higher simulated $C/N_{0}$ values, causing the estimated value to negatively diverge from the true one. A similar behavior is also observed at lower $C/N_{0}$ levels where estimated values become quickly unreliable as $C/N_{0}$ drops below 25 dB-Hz. Taking into account the chosen acquisition parameters and the simulation setup, the estimated $C/N_{0}$ range between 25 and approximately 40 dB-Hz are assumed reliable. Given the limited number of non-coherent integrations and the length of the coherent integration times adopted in this work, the $C/N_{0}$ values that can be observed considering the receiver sensitivity are expected to fall within this range in most of the tested scenarios.
@@
Figure 4: Comparison between expected and measured $C/N_{0}$. Generated $C/N_{0}$ (orange), estimated $C/N_{0}$ (black). Coherent integration time $T_{coh}:\ 4\ ms$, non coherent integrations: K: 15K:\ 15.
@@ 3.4 Extended Space Service Volume Simulator
The extended SSV simulator used in this work models the GNSS, RNSS, and SBAS constellations relevant to the analysis. It propagates broadcast ephemerides to determine satellite positions and generates the corresponding observables, including $C/N_{0}$, geometric Doppler shift, and pseudoranges. To estimate satellite visibility and received $C/N_{0}$ for L1 C/A and L5 signals at the spacecraft location, the simulator assigns to each satellite an antenna radiation pattern derived from the available literature. Table 4 summarizes the pattern adopted for each constellation.
High-fidelity three-dimensional gain patterns are available for GPS, Galileo, and QZSS on both L1 and L5; these are directly incorporated into the simulation. For BeiDou, complete antenna patterns have not yet been published at the time of writing. The simulator therefore employs the Min-Model, which provides a conservative representation of the main lobe for each band [31] and reconstructs the corresponding pattern through interpolation of the published values.
For other RNSS systems, detailed radiation patterns are similarly unavailable. In the case of GPS-Aided GEO Augmented Navigation (GAGAN) and IRNSS, an approximate antenna pattern is derived from CanX-2 mission measurements reported in [8]. This pattern is constructed by interpolating the available points conservatively, preserving the main lobe while assigning low gain at large off-boresight angles. For the remaining SBAS constellations, except for GAGAN, the BeiDou Min-Model is used for Geostationary Earth Orbit (GEO) and Inclined Geosynchronous Orbit (IGSO) satellites.
@@
For the remaining GNSS, RNSS, and SBAS constellations, calibration relies exclusively on the $C/N_{0}$ estimates obtained from the IQS snapshots collected during SC operations. Although this provides a less accurate calibration, due to the limited number of available measurements and the absence of detailed antenna patterns, it enables, for the first time, a simulation of multi-constellation signal availability in the cis-lunar environment based directly on experimental observations acquired in cis-lunar space and at the lunar surface. It is important to note that the power distribution among the various signal components for GPS, as well as for other constellations, may vary over time, for instance, in the case of GPS, due to the feature called flex power mode [24].
@@
GPS Official antenna pattern pre-flight measurements released by Lockheed Martin and Boeing. [27]
Galileo GRAP model. [12]
BeiDou (BDS-3) Min-model of the antenna main lobe, available per band and orbit type. [31] [11]
QZSS Official antenna pattern released by the Japan Aerospace Exploration Agency (JAXA). [6]
NavIC (IRNSS) Pattern available only for a submodule of the antenna. The patterns used correspond to BeiDou BDS-3 (B1) GEO-IGSO. [7]
@@
** Except for GAGAN [8].